OpenStudy (anonymous):

where does the graph of y=2x2-x-3 cross as the x-axis?

OpenStudy (shayaan_mustafa):

Hi ivogel3 :) How are you dear ? :D

OpenStudy (anonymous):

good

OpenStudy (shayaan_mustafa):

So we have to find the point on x-axis, known as x-intercept also. Right?

OpenStudy (anonymous):

yes i need to answers

OpenStudy (shayaan_mustafa):

Well we know at x-axis the coordinate of y is 0 And at y-axis the coordinate of x is 0. OK dear.

OpenStudy (anonymous):

ok

OpenStudy (shayaan_mustafa):

So in order to find x-intercept we have to put y=0

OpenStudy (anonymous):

if the equation is already equal to zero then what

OpenStudy (shayaan_mustafa):

Your equation becomes. \[\large x^2-x-3=0\] Now solve it.

OpenStudy (shayaan_mustafa):

Then just solve it simply.

OpenStudy (anonymous):

(x )(x ) but is other number that goes into the parenthasses

OpenStudy (shayaan_mustafa):

I know you can solve it now. You must get 2 answers.

OpenStudy (anonymous):

what about the 2

OpenStudy (zepp):

Do you know how to use the quadratic formula?

OpenStudy (zepp):

\(\huge x= \frac{-b \pm \sqrt{b^2-4ac}}{2a}\) Where \(a=2;~~b=-1;~~c=-3\)

OpenStudy (zepp):

\(\huge x= \frac{-(-1) \pm \sqrt{(-1)^2-4(2)(-3)}}{2(2)}\)

OpenStudy (zepp):

\(\huge x= \frac{1 \pm \sqrt{1+24}}{4}=\frac{1\pm\sqrt{25}}{4}=\frac{1\pm5}{4}\)

OpenStudy (anonymous):

X= -1 and x = 1.5

OpenStudy (zepp):

@dq101, please read the Code of Conduct; http://openstudy.com/code-of-conduct Don't give out straight answers :)

OpenStudy (anonymous):

@zepp, my bad! :$ I'm still new here!

OpenStudy (precal):